OpenStudy (anonymous):

simplify the expression: 2^2 * 2^3

OpenStudy (anonymous):

\[2^{2}\times 2^{3}\]

jimthompson5910 (jim_thompson5910):

the bases are the same so because you're multiplying the expressions, you add the exponents to get the final answer

jimthompson5910 (jim_thompson5910):

ex: 3^5 * 3^2 = 3^(5+2) = 3^7

OpenStudy (anonymous):

ok I wasn't sure if I multiplied those as well. thanks :)

jimthompson5910 (jim_thompson5910):

no you only multiply if you have something like (4^5)^2 ie you're raising an exponent to another exponent

OpenStudy (anonymous):

ok thanks again :)